# Exploring Parent’s Needs for Children-Centered AI to Support Preschoolers’ Storytelling and Reading Activities

### v1

@@ 3. METHOD
This paper aims to gain insights into parents’ practical adoption, usage, experiences, and perceptions toward AI-based intelligent storytelling technologies and tools for preschoolers between the ages of 3 and 6. We are particularly interested in investigating 1) current acceptance and usage of AI-based storytelling technologies and tools by parents with preschoolers, 2) their perceived benefits, 3) encountered challenges in using these tools, and 4) their perceptions and expectations of future AI-based intelligent storytelling technologies. To this end, we conducted a qualitative interview study with 17 parents. The rest of this section details the participants and recruitment, the types of data being collected, and the data analysis process.
@@ 3.1. Participants and Recruitment
We looked for participants who 1) were parents with preschoolers between the ages of 3 and 6 and 2) had previous or ongoing experience using AI-based intelligent tools to support storytelling activities. We did not specify what AI tools should be in our opinion and what kinds of AI tools participants had used. We recruited ones as long as they considered they had used storytelling tools with AI features for their storytelling practices with preschoolers. We also did not specify how much experience they should have, so that the recruitment could be open to people with various levels of usage and engagement in various kinds of AI-based storytelling technologies.
The participants were recruited via a snowball sampling approach. The first author, a native Mandarin speaker, contacted her classmates and friends who had preschoolers. After confirming that s/he had used at least one kind of AI-based storytelling tool, s/he then disclosed our research intention and recruited her/him for our interviews. Seven were contacted, and all agreed to be interviewed. To avoid bias from our own social networks, we then sought more participants by posting recruitment messages to online forums and communities. We designed the recruitment poster including self-introduction, research intention, recruitment criteria, and contact information of the first author, and team members distributed the poster to their social networks. Twelve reached out directly, and 10 of them agreed to be interviewed.
@@
P1
Mother
4-yo son, 10-yo daughter
Finance, Econ.
Company Staff
Interactive Talking Pen (C1), Robot (C5), Little Genius (C3), XiaoDu (C6)
@@
[0.5pt/0.5pt] P2
Mother
6-yo daughter, 12-yo son
Comp. Sci.
High School Staff
HuoHuoTu (C2), iPad (C4), Xiaodu (C6), Tmall Genie (C6)
@@
[0.5pt/0.5pt] P3
Father
5-yo daughter
Health Care
Company Employee
Talking Pen (C1), HuoHuoTu (C2), Early Learning Device (C3), iPad (C4), XiaoDu (C6)
@@
[0.5pt/0.5pt] P4
Father
5-yo daughter
Business Admin.
Company Manager
Jiaojiao (iPad app) (C4), Qiaohu (an educational brand, including TV programs, animated series, interactive games, and educational materials)
@@
[0.5pt/0.5pt] P5
Mother
6-yo daughter
Automation
CIO
Interactive Pen (C1), Interactive Player (C2), iPad (C4), Robot (C5)
@@
[0.5pt/0.5pt] P6
Mother
6-yo daughter
Business Admin.
Data Company Employee
HuoHuoTu (C2), Smartwatch with story reading apps, iPad (C4)
@@
[0.5pt/0.5pt] P7
Mother
5-yo son
Educational Psychology
University Staff
Robot (C5)
@@
[0.5pt/0.5pt] P8
Mother
6-yo son
Comp. Sci.
Researcher
Interactive Talking Pen (C1), Seewo (C3), Luka (C5), iPad (C4)
@@
[0.5pt/0.5pt] P9
Mother
6-yo son
Finance, Econ.
University Financial Staff
Interactive Talking Pen (C1)
@@
[0.5pt/0.5pt] P10
Mother
5-yo daughter
Public Admin.
Company Employee
Qiaohu
@@
[0.5pt/0.5pt] P11
Father
4-yo daughter
Comp. Sci.
University Researcher
Interactive Pen (C1), Early Learning Device (C3), Robot (C5), Siri (C6), Qiaohu
@@
[0.5pt/0.5pt] P12
Father
6-yo daughter
Education
University Staff
Tmall Genie (C6), XiaoAi (C6), iPad (C4)
@@
[0.5pt/0.5pt] P13
Father
3-yo son, 5-yo daughter
Public Admin.
AI4EDU Institute Staff
MiXiaoTu (C2), iPad (C4)
@@
[0.5pt/0.5pt] P14
Mother
3-yo daughter, 7-yo son
Comp. Sci.
Company Employee
HuoHuoTu (C2), Alpha Egg (C5), Tmall Genie (C6), XiaoAi (C6)
@@
[0.5pt/0.5pt] P15
Father
3-yo daughter, 10-yo son
Psychotechnics
AI4EDU Institute Researcher
Interactive Talking Pen (C1), Storyplayer (C2), Alpha Egg (C5), Tmall Genie (C6)
@@
[0.5pt/0.5pt] P16
Mother
4-yo son
Lawyer
Company Employee
Interactive Talking Pen (C1), Qiaohu, Tmall Genie (C6)
@@
[0.5pt/0.5pt] P17
Father
3-yo daughter, 7-yo daughter
Education
Kindergartener
Tmall Genie (C6)
@@ 3.2. Data Collection
The interviews were semi-structured, including questions about 1) their daily storytelling practices and experiences, 2) the AI-based intelligent storytelling tools they have used and how they experienced these tools (when the technologies and tools reported by participants did not align well with research communities’ universal definition of AI technologies, we explored as much as possible the ”intelligent functions” that parents considered from these technologies and tools), 3) the benefits and challenges they experienced, and 4) their expectations of future intelligent tools. For parents who have used LLM-based tools (e.g. ChatGPT), we specifically asked 5) how they considered if LLM-based technologies were used to support preschooler-oriented storytelling. All interview questions were designed to be general so that the questions could be inclusive to participants to share various levels of experiences with different technology-mediated storytelling practices.
The interviews were mainly conducted in Mandarin by the first author. 12 were conducted via WeChat Voice call and Tencent Meeting, and 5 were conducted in person. Each interview lasted for 40-60 minutes, with a 100 CNY honorarium. During the interview, we encouraged participants to provide detailed demonstrations of how these tools functioned and how they used them. When some interesting points or prior experiences were mentioned, we asked appropriate follow-up questions for more details and concrete examples. With participants’ permission, we also obtained other forms of data to triangulate their interview responses. These data included 1) screenshots of mobile applications the participants have used and 2) related photos or videos they shot, if any.
@@ 3.3. Data Analysis
With the participants’ permission, we recorded the audio of all interviews and transcribed them into Chinese verbatim after the interviews for analysis. We followed thematic analysis as outlined by Clarke et al. (2015) to analyze the main themes in the interview data in an inductive way (Thomas, 2006). Three authors participated in the data analysis process. We started coding and analysis while the data was being collected. We iterated through generating codes and themes from collected data and checking and elaborating these codes by collecting more data. The data collection phase ended when recurring themes were identified.
@@
During the open coding phase, three authors independently reviewed the collected data and generated codes pertaining to our research questions. We had regular meetings on a weekly basis to discuss codes and ensure reliability. After this stage, we generated the initial code list, capturing participants’ practices, motivations, and perceptions related to their daily storytelling actions; their usage, motivations, and experiences related to the intelligent storytelling tools they had used, as well as their perceptions and expectations to the general AI-based storytelling tools.
@@
Based on this code list, we then re-focused our analysis at the broader level of themes, and synthesized and grouped the codes to create higher-level overarching themes. The initial analysis centered on themes related to storytelling tools parents had used, parents’ adoption, usage and experiences of these tools, the difficulties and challenges they have encountered, their children’s adoption, usage and experiences of these tools, parents’ perceptions and concerns about LLM-based AI tools, and parents’ expectations of future AI technologies. As analysis proceeded, we found that, due to our open criteria to “what AI tools should be in our opinion and what kinds of AI tools participants had used”, some tools reported by our participants, such as Huohuotu, story player etc., were not aligned well with research communities’ typical definition of AI technologies [^2] https://en.wikipedia.org/wiki/Artificial_intelligence on the one hand, and some, such as part of apps in iPad, were not used for story reading and telling, but for learning on the other hand. Considering the research aim of “gaining insights into parents’ practical adoption, use, experiences, and perceptions toward AI-based storytelling technologies and tools, through which to shed light on how we can better design future AI tools to support parent-preschooler interactive storytelling”, we removed codes regarding parents’ practical usage and experience of these not so AI-based tools and not storytelling-oriented tools to focus the analysis more on the content relevant to our research questions. In addition, we realized that the perceptions of using existing intelligent storytelling tools and LLM-based AI tools were closely related. We then further focused on themes that reflect similarities and differences between existing intelligent storytelling tools and LLM-based AI tools.
@@
Through several rounds of analysis and discussions, three authors reached an agreement and generated our final satisfactory thematic map of the data, which included four primary themes: 1) parents’ attitudes and perception to AI-based storytelling with preschoolers, 2) their perceived benefits, and encounter 3) interactive and 4) algorithmic challenges in practice. In the following section, we present the full details of these themes. Throughout, we use representative quotes from our participants, translated from Chinese into English by the first author, to illustrate our points. To protect our participants’ identities, we use P1, P2, etc., to denote our study participants.
@@ 3.4. Positionality of the Authors
We acknowledge that, for qualitative research, the research perspectives and approaches are shaped by our own experiences and positionality (Pascoe, 2022). Specifically, we are researchers living and working in the U.S. and China, primarily researching human-centered AI and the perceived trustworthiness, fairness and effectiveness of AI in education. We come from a mix of disciplinary backgrounds, including Computer Science, Learning Sciences and Technology, Education, and HCI, which we have drawn on to conduct prior research into sociotechnical approaches to human-centered AI design practices. We have experience working with school teachers and district personnel on technology integration, and AI practitioners on projects related to human-centered educational technology design. In general, we recognize the tremendous potential of the rapidly developed AI in promoting societal progress and enhancing people’s quality of life and work. However, we do not agree with the notion that AI will replace human roles, such as the role of parents in children’s storytelling. We firmly believe that the optimal design approach should be human-centered, offering greater flexibility and adaptability. Our results confirm the validity of our hypothesis.
@@ 3.5. Ethical Considerations
All researchers received formal ethics training, and were granted official certificates from the ethics committee in their institute. During the study, we took careful steps to protect user privacy. Before we started interviews, we informed participants of our intention and background information, and got permission from them. All data collected during our study was used in an anonymized way, i.e., there was no link between the collected data and an individual user. We strictly abide by the confidentiality agreement, which promises the dataset is only used for this research and any form of information leakage is not permitted. Our institute does not fall under the jurisdiction of an internal review board or similar ethics committee. We did, however, get feedback from peers inside and outside our faculty to validate the ethical perspective of our research.
@@ 4. Findings
Our findings generally suggested that it had been very common among parents to utilize various types of AI-based technologies and tools to facilitate their daily parent-children storytelling activities (as shown in the right column of Table 1). Among 17 parents we interviewed, 14 had used more than two types of technologies to support their daily parent-children storytelling practices. According to the interaction modes and functionalities, we roughly classified these tools into six categories, using C1-C6 to present. Specific details can be found in Table 2. Some of these tools were aligned well with research communities’ typical definition of AI technologies, while some others, such as tools of C1 and C2, were not. In the following, we will elaborate on the experiences, perceived benefits, and encountered challenges of our participants in using these tools.
4.1. Parents’ Attitudes and Expectations to AI-Based Storytelling Technologies
From parents’ perspective, our participants expressed mixed (positive and negative) attitudes towards existing AI-based storytelling technology. We found these attitudes were somewhat correlated with parents’ educational background, work experience, and parenting philosophies. Regardless of their attitudes, they all believed AI’s potential to support storytelling scenarios for preschoolers. However, they emphasized the functionalities it possessed and the role it played required careful attention. All of our participants emphasized that no matter how intelligent AI technologies become, they should be parents’ assistants in storytelling practices, rather than replacing parents to work. In the following sections, we will elaborate on parents’ attitudes and expectations regarding AI-based storytelling tools.
4.1.1. Parents’ Attitudes to AI-based Storytelling for Preschoolers
Our participants’ attitudes towards existing AI-based storytelling technologies varied from individual to individual. Some were generally positive and while others were relatively negative. Further, we found parents’ individual educational or occupational backgrounds influenced their individualized adoption and attitudes to AI-based storytelling tools to some extent. Among our interviewees, five participants were engaged in technology-related professions, or with technology-related educational background (P5 was a CIO of a company, P6 worked in the field of digital healthcare, P8 and P11 were involved in computational science research, and P13 held a managerial position in an intelligent education research institution). In contrast to other participants, these five interviewees exhibited more positive attitudes and higher expectations for future AI-based storytelling technologies. They believed that despite the current limitations of AI-based storytelling tools, there was great potential for future AI technologies to address these limitations and better assist parent-children storytelling. They expressed their eager anticipation to the emergence of such technologies. In contrast, four participants with educational backgrounds or occupations in the educational field (P2, P7, P12 and P17), hold relatively conservative attitudes toward AI-based storytelling tools. During the interviews with them, they explicitly expressed strong dislike and even disapproval of preschoolers’ use of AI-based tools, with the reasons of challenge between preschoolers’ limited discriminatory abilities and almost uncontrollable AI (P2), challenge between preschoolers’ cognitive development level and AI’s virtual interaction mode (P7), challenge between preschooler’ free, personalized development and AI’s overly direct goal orientation and standardization (P12) etc.
Meanwhile, parents’ parenting philosophies also significantly impact their attitudes toward accepting and using AI-based storytelling technologies. Throughout our interviews, participants often articulated their viewpoints using statements like “I believe that education for children should be XX; in this philosophy, AI XX.” Our data indicated that parenting philosophies had a more direct impact on parents’ acceptance and use of AI tools than educational or occupational backgrounds. For example, despite P2 having a background in computer science and P15 working in an intelligent education research institution, their attitudes toward AI tools are not particularly positive. Instead, they lean towards the idea of parents personally accompanying their children in storytelling, rather than relying on a smart machine.
4.1.2. Can AI Replace Parents’ Roles in Storytelling?
@@
None of the participants believed AI technologies could replace the role of parents in storytelling scenarios because “it cannot support interactions with emotion and warmth, and cannot replace parents’ love, care and hugs” [P13, a father with a 3-year-old son and a 5-year-old daughter]. They emphasized that story reading and telling with preschoolers was a very children-centered and comprehensive process. Through the storytelling process, they hoped children could perceive the world, learn different knowledge and skills, and feel different cultural contexts. At the same time, they also pointed out that it was not only for learning and education, but also for guiding children’s active thinking, cultivating their reading habits, or just be there and spend parent-child time with my child [P5].
Yet, AI-based storytelling tools paid more attention to reading the story or teaching knowledge. They couldn’t be able to care for and love them as their parents did. Even the more powerful LLM-based applications couldn’t replace parents to carry out such a children-centered comprehensive process. P9, a mother with a 6-year-old son strongly opposed the existing trend of using AI technology to tell stories for preschoolers. She believed that children’s growth should be accompanied by parents, rather than a smart machine.
@@
“You propose to design a super smart machine and ask me to use it to educate my son. Why? I don’t understand, I think the most important thing is the interaction between parents and children, between human and human. Mothers should not find a machine to accompany their children, but rather ‘I accompany you’. A machine is just a machine, no matter how advanced it is, it is still just a machine.”[P9]
@@
At the same time, many of our participants thought that AI-mediated storytelling was not real storytelling or reading. They emphasized that storytelling and reading in the real world was not just about telling and reading those stories, asking the questions, or teaching some knowledge, but also the reading habits, thinking methods, and even postures and attitudes of reading. Parents need to pay attention to all of these aspects, but AI tools often only read the stories.
While all our participants agreed that AI technologies absolutely couldn’t replace the role of parents in practical storytelling scenarios, most of them explicitly expressed their expectations of intelligent technologies to become their “assistants” or “secret weapons”, with functions of freeing up parents’ physical labor (e.g. stroybook reading), helping them generate valuable and high-quality interactive content or questions, or inspiring them on how to interact with their children.

### v2

@@ 3. METHOD
This paper aims to gain insights into parents’ practical adoption, usage, experiences, and perceptions toward AI-based intelligent storytelling technologies and tools for preschoolers between the ages of 3 and 6. We are particularly interested in investigating 1) current acceptance and usage of AI-based storytelling technologies and tools by parents with preschoolers, 2) their perceived benefits, 3) encountered challenges in using these tools, and 4) their perceptions and expectations of future AI-based intelligent storytelling technologies. To this end, we conducted a qualitative interview study with 17 parents. We iterated through generating codes from collected data and checking and elaborating these codes by collecting more data. We stopped data collection once all the core variables were saturated. The rest of this section details the participants and recruitment, the types of data being collected, and the data analysis process.
@@ 3.1. Participants and Recruitment
We looked for participants who 1) were parents with preschoolers between the ages of 3 and 6 and 2) had previous or ongoing experience using AI-based intelligent tools to support storytelling activities. We did not specify what AI tools should be in our opinion, and what kinds of AI tools participants had used. We recruited those as long as they considered that they had used storytelling tools with AI features for their storytelling practices with preschoolers. We also did not specify how much experience they should have, so that the recruitment could be open to people with various levels of usage and engagement in various kinds of AI-based storytelling technologies.
The participants were recruited via a snowball sampling approach. The first author, a native Mandarin speaker, contacted her classmates and friends who had preschoolers. After confirming that s/he had used at least one kind of AI-based storytelling tool, the first author then disclosed our research intention and recruited her/him for our interviews. Seven were contacted, and all agreed to be interviewed. To avoid bias from our own social networks, we also sought participants by posting recruitment messages on online forums and communities. We designed the recruitment poster including self-introduction, research intention, recruitment criteria, and contact information of the first author, and team members distributed the poster on their social networks. Twelve reached out directly, and ten of them agreed to be interviewed.
@@
P1 Mother 4-yo son, 10-yo daughter Finance, Econ. Company Staff Interactive Talking Pen (C1), Robot (C5), Little Genius (C3), XiaoDu (C6)
@@ 3.2. Data Collection
The interviews were semi-structured, including questions about 1) their daily storytelling practices and experiences, 2) the AI-based intelligent storytelling tools they have used and how they experienced these tools (when the technologies and tools reported by participants did not align well with research communities’ definition of AI technologies, we explored as much as possible the “intelligent functions” that parents considered from these technologies and tools) and prompted them by illustrating typical “AI technologies”, 3) the benefits and challenges they experienced, and 4) their expectations of future intelligent tools. For parents who have used LLM-based tools (e.g. ChatGPT), we specifically asked 5) how they considered if LLM-based technologies were used to support preschooler-oriented storytelling. All interview questions were designed to be general so that the questions could be inclusive to participants to share various levels of experiences with different technology-mediated storytelling practices.
The interviews were mainly conducted in Mandarin by the first author. 12 were conducted via WeChat Voice call and Tencent Meeting, and 5 were conducted in person. Each interview lasted for 40-60 minutes, with a 100 CNY honorarium. During the interview, we encouraged participants to provide detailed demonstrations of how these tools functioned and how they used them. When some interesting points or prior experiences were mentioned, we asked follow-up questions for more details and concrete examples. With participants’ permission, we also obtained other forms of data to triangulate their interview responses. These data included 1) screenshots of mobile applications the participants have used and 2) related photos or videos they took, if any.
@@ 3.3. Data Analysis
With the participants’ permission, we recorded the audio of all interviews and transcribed them into Chinese verbatim after the interviews for analysis. We followed thematic analysis as outlined by Clarke et al. (2015) to analyze the main themes in the interview data in an inductive way (Thomas, 2006). Three authors participated in the data analysis process. We started analysis while the data was being collected. During the open coding phase, three authors independently reviewed the collected data and generated codes pertaining to our research questions. We had regular meetings on a weekly basis to discuss codes and ensure reliability. After this stage, we generated the initial code list, capturing participants’ practices, motivations, and perceptions related to their daily storytelling actions; their usage, motivations, and experiences related to the intelligent storytelling tools they had used, as well as their perceptions and expectations to the general AI-based storytelling tools.
Based on this code list, we then re-focused our analysis at the broader level of themes, and synthesized and grouped the codes to create higher-level overarching themes. The initial analysis centered on themes related to storytelling tools parents had used, parents’ adoption, usage and experiences of these tools, the difficulties and challenges they have encountered, their children’s adoption, usage and experiences of these tools, parents’ perceptions and concerns about LLM-based AI tools, and parents’ expectations of future AI technologies.
@@
As analysis proceeded, we found that, due to our open criteria to “what AI tools should be and what kinds of AI tools participants had used”, some tools reported by our participants, such as Huohuotu, story player, etc., were not aligned well with research communities’ typical definition of AI technologies (referring to technologies that enable computers and machines to learn, read, write, create and analyze, simulating human intelligence and problem-solving capabilities [^2] https://www.ibm.com/topics/artificialintelligence ) on the one hand, and some, such as part of applications in iPad, were not used for story reading and telling, but for learning on the other hand. Considering the research aim of “gaining insights into parents’ practical adoption, use, experiences, and perceptions toward AI-based storytelling technologies and tools, through which to shed light on how we can better design future AI tools to support parent-preschooler interactive storytelling”, we removed codes regarding parents’ practical usage code of these tools to focus the analysis more on the codes relevant to our research questions. We still retained data on participants’ experiences with these technologies because we considered that these data expressed parents’ actual needs for storytelling technologies. In addition, we realized that the perceptions of using existing intelligent storytelling tools and LLM-based AI tools were closely related. We then further focused on themes that reflect similarities and differences between existing intelligent storytelling tools and LLM-based AI tools.
@@
Through several rounds of analysis and discussions, three authors reached an agreement and generated our final satisfactory thematic map of the data, which included four primary themes: 1) parents’ attitudes and perception of AI-based storytelling with preschoolers, 2) their perceived benefits, and encountered 3) interactive and 4) algorithmic challenges in practice. In the following section, we present the full details of these themes. Throughout, we use representative quotes from our participants, translated from Chinese into English by the first author, to illustrate our points. To protect our participants’ identities, we use P1, P2, etc., to denote our study participants.
@@ 3.4. Positionality of the Authors
We acknowledge that, for qualitative research, the research perspectives and approaches are shaped by our own experiences and positionality (Pascoe, 2022). Specifically, we are researchers living and working in the China and U.S., primarily researching human-centered AI and the perceived trustworthiness, fairness and effectiveness of AI in education. We come from a mix of disciplinary backgrounds, including Computer Science, Learning Sciences and Technology, Education, and HCI, which we have drawn on to conduct prior research into sociotechnical approaches to human-centered AI design practices. We have experience working with school teachers and district personnel on technology integration, and AI practitioners on projects related to human-centered educational technology design. In general, we recognize the tremendous potential of the rapidly developed AI in promoting societal progress and enhancing people’s quality of life and work. However, we disagree with the notion that AI will replace human roles, such as the role of parents in children’s storytelling. We firmly believe that the optimal design approach should be human-centered, offering greater flexibility and adaptability. Our results confirm the validity of our hypothesis.
@@ 3.5. Ethical Considerations
All researchers received formal ethics training, and were granted official certificates from the ethics committee in their institute. During the study, we took careful steps to protect user privacy. Before we started interviews, we informed participants of our intention and background information, and got permission from them. All data collected during our study was used in an anonymized way, i.e., there was no link between the collected data and an individual user. We strictly abide by the confidentiality agreement, which promises the collected data is only used for this research and any form of information leakage is not permitted.
@@ 4. Findings
Our findings suggested that it had been very common among our participants to utilize various types of AI-based technologies and tools to facilitate their daily parent-children storytelling activities (as shown in the right column of Table 1). Among 17 parents we interviewed, 14 had used more than two types of technologies to support their daily parent-children storytelling practices. According to the interaction modes and functionalities, we roughly classified these tools into six categories. Specific details can be found in Table 2. Some of these tools were aligned well with research communities’ typical definition of AI technologies, while some others, such as tools of C1 and C2, were not. In the following, we will elaborate on the experiences, perceived benefits, and encountered challenges of our participants when using these tools.
4.1. Parents’ Attitudes and Expectations to AI-Based Interactive Storytelling
4.1.1. Parents’ Attitudes to AI-based Storytelling Tools: Cautiously Positive
Generally, our participants expressed a cautiously positive attitude toward existing AI-based storytelling tools. That is, although they believed AI’s potential in supporting interactive storytelling for preschoolers and reported some benefits of existing tools they perceived (see section 4.2 for details), they emphasized a series of concerns to and limitations of these tools (see section 4.3 and 4.4 for details).
@@ 4.1.1. Parents’ Attitudes to AI-based Storytelling Tools: Cautiously Positive
Further, none of the participants believed AI technologies, no matter how they became, could replace the role of parents in interactive storytelling scenarios, because “it cannot support interactions with emotion and warmth, and cannot replace parents’ love, care and hugs[P13, a father with a 3-year-old son and a 5-year-old daughter]. They emphasized that interactive storytelling with preschoolers was a very child-centered comprehensive process. Through this process, they hoped children could perceive the world, learn different knowledge and skills, and feel different cultural contexts, and hoped to guide children’s active thinking, cultivate their reading habits, or just be there and spend parent-child time with my child [P5]. Yet, AI-based tools only paid attention to reading the story or teaching knowledge. P9, a mother with a 6-year-old son strongly opposed the existing trend of using AI technologies to read stories for preschoolers. She believed that children’s growth should be accompanied by parents, rather than a smart machine.
@@
“You propose to design a super smart machine and ask me to use it to educate my son. Why? I don’t understand. I think the most important thing is the interaction between parent and child, between human and human. Mothers should not find a machine to accompany their children, but rather ‘I accompany you’. A machine is just a machine, no matter how advanced it is, it is still just a machine.”[P9]
@@
In addition, many of our participants thought that AI-based interactive storytelling was not real storytelling and reading. They emphasized that storytelling and reading in the real world was not just about reading stories, asking questions, or teaching some knowledge, but also the reading habits, thinking methods, and even postures and attitudes of reading. Parents need to pay attention to all of these aspects, but AI tools only read the stories and teach knowledge.
Significantly, our findings suggested that parents’ attitudes were somewhat correlated with their educational background, work experience, and parenting philosophies. Specifically, we found parents’ educational and occupational backgrounds influenced their adoption and attitudes to AI-based storytelling tools to some extent. Among our participants, five were engaged in technology-related professions, or with technology-related educational backgrounds (P5 was a CIO of a company, P6 worked in the field of digital healthcare, P8 and P11 were involved in computational science research, and P13 held a managerial position in an intelligent education research institution). In contrast to the other participants, these five participants exhibited more positive attitudes and higher expectations for AI-based interactive storytelling technologies. They believed that despite the current limitations of AI technologies, there was great potential for future technologies to address these limitations and better assist parent-children interactive storytelling. They expressed their eager anticipation of the emergence of such technologies. In contrast, four participants with educational backgrounds or occupations in the educational field (P2, P7, P12, and P17), hold relatively conservative attitudes toward AI-based interactive storytelling tools. During the interviews with them, they explicitly expressed strong dislike and even disapproval of preschoolers’ use of AI-based tools, with the reasons of tentions between preschoolers’ limited discriminatory abilities and uncontrollable AI (P2), between preschoolers’ cognitive development level and AI’s virtual interaction modes (P7), between preschooler’ free, personalized development and AI’s overly direct goal orientation and standardization (P12), etc.
Parenting philosophies also significantly impact our participants’ attitudes toward AI-based storytelling technologies. Throughout our interviews, participants often articulated their viewpoints using statements like “I believe that education for children should be XX; in this philosophy, AI should XX.” Our data indicated that parenting philosophies had a more direct impact on parents’ acceptance and use of AI tools than educational or occupational backgrounds. For example, despite P2 having a background in computer science and P15 working in an intelligent educational research institution, their attitudes toward AI tools were not particularly positive. Instead, they lean towards the idea of parents personally accompanying their children in storytelling, rather than relying on a smart machine.
4.1.2. Parents’ Expectations to AI-based Storytelling Tools: “Being Parents’ Secret Weapon
While all our participants agreed that AI technologies absolutely couldn’t replace parents’ role in practical interactive storytelling scenarios, they expected AI-based tools to be their ‘secret weapon’ to help free up their physical labor, assist in meaningful parent-child interaction, and help better understand children and find resources according to their personalized characteristics and needs.
Freeing up physical labor. Many of our participants expressed that telling stories to children was a very tiring job, and it became even more difficult if interaction was required. They therefore expected a machine could help them read stories, then they could focus on more valuable interactions with children. P14, a mother of a 7-year-old son and a 3-year-old daughter, expressed her strong expectation of technology-supported story reading tools: “Reading for 20 minutes every day often makes me exhausted. It is a kind of physical exhaustion of the mouth. Interactive storytelling is more tiring. I’m often tired when I come home from work. Sometimes, when I tell my child a story, I end up falling asleep myself. If a robot can read the story, it will be very great.
Assisting meaningful parent-child interaction. Given the practical challenges many parents faced in conducting meaningful and high-quality interactions with their children, our participants hoped AI tools could assist their parent-child interaction, such as generating valuable interactive content, and inspiring parents on how to interact with children. As P9 said, “If my son asked me a question and I didn’t know the answer, I would tell him, ‘Wait a moment, let me ask my secret weapon’. After finding the answer, I would tell him. I thought this was a very cool thing.” She emphasized that the important part was that the secret weapon told the answer to parents, and helped parents to better interact with children, rather than the machine directly interacting with children.
Helping to better understand children and find resources according to their personalized characteristics and needs. With the explosive growth of digital learning resources, providing valuable and suitable learning materials for children at specific stages was a common challenge faced by parents. They consequently hoped AI technologies could help them better understand their children and find valuable and suitable resources according to their personalized characteristics and needs. Many parents in our study expressed that they were not satisfied with many currently used tools, but they still used them mainly because “I don’t have to find resources by myself.”